# Families of cyclic curve coverings with maximal monodromy

### v1

@@ 5. The Coleman-Oort conjecture
and, by definition, $\mathsf{Z}$ is special if and only if $\mathsf{Z}=S_{f}$. Proving the necessity of condition ($\star$5.1) would thus amount to showing that $\mathsf{Z}$ is special if and only if $\mathsf{Z}=S_{f}=S(G)$.
@@
To complete the picture, we recall that special subvarieties are subvarieties in $\mathcal{A}_{g}$, which are linear with respect to the locally symmetric metric induced from the Siegel space $\mathfrak{S}_{g}$. More precisely, denote by $\pi:\mathfrak{S}_{g}\rightarrow A_{g}$ the natural projection map and assume that an algebraic subvariety $Z\subset\mathcal{A}_{g}$ is totally geodesic if $Z=\pi(X)$, for a totally geodesic submanifold $X\subset\mathfrak{S}_{g}$. Then, special subvarieties are totally geodesic (and not all totally geodesic subvariety of $\mathcal{A}_{g}$ are special). Thus, if we denote by $X_{f}$ the smallest totally geodesic subvariety of $\mathcal{A}_{g}$ containing our family $\mathsf{Z}$, we have:
@@
In this setting, our Theorem 4.7, reads as follows.
@@
(5.1) $\operatorname{Mon}^{0}(JC_{b})=\operatorname{Mon}^{0}(JC^{\prime}_{b})\times\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))=\operatorname{Sp}_{2}\times\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
@@
By Lemma 4.6, the covering $D^{\nu}\rightarrow D^{\nu}$ has still no repeating factions in the monodromy. Furthermore, by Theorem 3.3, we have $\operatorname{Prym}(\tilde{D},D)\sim\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu})$. Also, recall that $g(D^{\nu})=0$ and hence $JD^{\nu}\sim\operatorname{Prym}(\tilde{D},D)$. Therefore, by Theorem 4.4
$\operatorname{Mon}^{0}(JD^{\nu})=\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))=\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
@@
The examples of special families of cyclic coverings can be found in [10, 11]. The proof of the statement is based in the following considerations. First, if the family is one-dimensional, then the uniformizing symmetric space of $S(G)$ is $\mathfrak{S}_{1}$, so this is automatic. Families with double and triple coverings always have no repeating factors in the monodromy; this is discussed in section 6. Finally, for the remaining higher dimensional families we refer the reader to [27, Theorem 1.2], where it is shown that their uniformizing symmetric space is an irreducible symmetric space, and this implies that only one factor appears in $(\operatorname{Sp}^{G})^{ad}$ (hence, in particular, they have no repeating factors in the monodromy). ∎
@@
$\operatorname{Mon}^{0}=(\operatorname{Sp}(H^{1}(C,{\mathbb{Q}}))^{G})^{der}$
as soon as $g(C^{\prime})\geq 4$. This implies
$\operatorname{Mon}^{0,ad}=(\operatorname{Sp}(H^{1}(C,{\mathbb{Q}}))^{G})^{ad},$
namely it guarantees the maximality of the monodromy in the sense of our Theorem 4.7. Hence, for $g^{\prime}\geq 4$ our result provides only an alternative technique to get the maximality in our hypothesis. On the other hand, our result also covers the genera $g^{\prime}\leq 4$ for cyclic coverings with no repeating factors in the monodromy. We also note that [17, Theorem 1.3] implies that $S_{f}=S(G)$ for every family of abelian covers with $g^{\prime}\geq 4$. Thus, arguing as in the proof of Theorem 5.5 and, using that by [13, Theorem 1.1] if $g^{\prime}\geq 2$ there are no families satisfying condition ($\star$5.1), then we can exclude the existence of families of abelian covers with $g^{\prime}\geq 4$ yielding special (or even totally geodesic) subvarieties generically contained in the Torelli locus.

### v2

@@ 5. The Coleman-Oort conjecture
and, by definition, $\mathsf{Z}$ is special if and only if $\mathsf{Z}=S_{f}$. Proving the necessity of condition ($\star$5.1) thus amount to showing that $\mathsf{Z}$ is special if and only if $\mathsf{Z}=S_{f}=S(G)$.
@@
To complete the picture, we recall that special subvarieties are subvarieties in $\mathcal{A}_{g}$, which are linear with respect to the locally symmetric metric induced from the Siegel space $\mathfrak{S}_{g}$. More precisely, denote by $\pi:\mathfrak{S}_{g}\rightarrow A_{g}$ the natural projection map. An algebraic subvariety $Z\subset\mathcal{A}_{g}$ is totally geodesic if $Z=\pi(X)$, for a totally geodesic submanifold $X\subset\mathfrak{S}_{g}$. Then, special subvarieties are totally geodesic (and not all totally geodesic subvariety of $\mathcal{A}_{g}$ are special). Thus, if we denote by $X_{f}$ the smallest totally geodesic subvariety of $\mathcal{A}_{g}$ containing our family $\mathsf{Z}$, we have:
@@
In this setting, our Theorem 4.7 implies:
@@
Proof.
The special subvariety $S(G)$ is the orbit of $\operatorname{Sp}^{G}$. This is the same as the orbit of $(\operatorname{Sp}^{G})^{ad}$. Since $(\operatorname{Sp}^{G})^{ad}=((\operatorname{Sp}^{G})^{der})^{ad}$, from (4.4), we have
$(\operatorname{Sp}^{G})^{ad}=\prod_{\chi\neq\bar{\chi}}\operatorname{PSU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{PSp}(2m_{\chi},\mathbb{R}).$
Let $Q$ be one of the simple factors and denote by $\delta(Q)$ the dimension of the corresponding symmetric space. Note that when $Q$ is compact we have $\delta(Q)=0$. Therefore,
$\dim S(G)$ $=\sum_{\chi}\delta(\operatorname{PSU}(m_{\chi},m_{\bar{\chi}}))+\sum_{\chi}\delta(\operatorname{PSp}(2m_{\chi},\mathbb{R}))$
$=\sum_{m_{\chi}m_{\bar{\chi}}\neq 0}\delta(\operatorname{PSU}(m_{\chi},m_{\bar{\chi}})+\sum_{m_{\chi}\neq 0}\delta(\operatorname{PSp}(2m_{\chi},\mathbb{R})).$
It follows that the dimension of the symmetric space obtained taking the orbit of $\operatorname{Sp}^{G}$ is the same as the one coming from $(\operatorname{Sp}^{G})^{pos}$. Since $X_{f}$ is the totally geodesic subvariety associated with $\operatorname{Mon}^{0}_{\mathbb{R}}$, Theorem 4.7 yields $X_{f}=S(G)$. Hence, the statement holds. ∎
@@
(5.1) $\operatorname{Mon}^{0}(JC_{b})$ $=\operatorname{Mon}^{0}(JC^{\prime}_{b})\times\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))$
$=\operatorname{Sp}_{2}\times\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
@@
By Lemma 4.6, the covering $D^{\nu}\rightarrow D^{\nu}$ has still no repeating factions in the monodromy. Furthermore, by Theorem 3.3, we have $\operatorname{Prym}(\tilde{D},D)\sim\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu})$. Also, recall that $g(D^{\nu})=0$ and hence $J\tilde{D}^{\nu}\sim\operatorname{Prym}(\tilde{D},D)$. Therefore, by Theorem 4.4
$\operatorname{Mon}^{0}(J\tilde{D}^{\nu})=\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))=\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
@@
The examples of special families of cyclic coverings can be found in [10, 11]. The proof of the statement is based in the following considerations. First, if the family is one-dimensional, then the uniformizing symmetric space of $S(G)$ is $\mathfrak{S}_{1}$, so this is automatic. Families with double and triple coverings always have no repeating factors in the monodromy; this is discussed in section 6. Finally, for the remaining higher dimensional families we refer the reader to [26, Theorem 1.2], where it is shown that their uniformizing symmetric space is an irreducible symmetric space, and this implies that only one factor appears in $(\operatorname{Sp}^{G})^{pos}$ (hence, in particular, they have no repeating factors in the monodromy). ∎
@@
$\operatorname{Mon}^{0}_{\mathbb{R}}=(\operatorname{Sp}(H^{1}(C,\mathbb{R}))^{G})^{der}$
as soon as $g(C^{\prime})\geq 4$. Hence, for $g^{\prime}\geq 4$ our result provides only an alternative technique to get the maximality in our hypothesis. On the other hand, our result also covers the genera $1\leq g^{\prime}\leq 4$ for cyclic coverings with no repeating factors in the monodromy. We also note that [17, Theorem 1.3] says that, letting $d$ be the maximal dimension of an irreducible complex representation of any finite group $G$, then
$X_{f}=S_{f}=S(G)$
for every family of unramified $G$-covers with $g^{\prime}\geq 2d+2$ or ramified $G$-covers with $g^{\prime}>max\{2d+1,d^{2}\}$. Thus, arguing as in the proof of Theorem 5.5 and, using that by [13, Theorem 1.1] if $g^{\prime}\geq 2$ there are no families satisfying condition ($\star$5.1), we get:
Proposition 5.10.
Let $G$ be any finite group and let $d$ be the maximal dimension of an irreducible complex $G$-representation. Then, every family of unramified $G$-covers with $g^{\prime}\geq 2d+2$ or ramified $G$-covers with $g^{\prime}>max\{2d+1,d^{2}\}$ is not totally geodesic.